Circular No. 8327

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

      URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 G2 (SOHO)

Further to IAUC 8309, K. Battams reports his measures for a comet (discovery observation below) found by T. Hoffman on C2 coronagraph images posted at the SOHO website. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2004-H01. The object appears to be a member of the Meyer group (cf. IAUC 7832).

2004 UT  α_{2000}  δ_{2000}
Apr. 10.064  1^{h}19.^{m}6  +10^{o}02′

COMETS C/2003 K10–K12, C/2003 L3–L8, C/2003 M1–M7 (SOHO)

Battams and K. McGleam also report their measurements of numerous Kreutz sungrazing comets found on SOHO C2 website images by P. Shkreby (C/2003 K10 and L6), D.-h. Chen (C/2003 K11), T. Hoffman (C/2003 K12), X. Leprette (C/2003 L3), R. Kracht (C/2003 L4, L5, L8, M1, M3, M4, and M6), M. Meyer (C/2003 L4), M. Boschat (C/2003 L7 and M5), K. Černis (C/2003 M1), J. Sachs (C/2003 M2), and X.-m. Zhou (C/2003 M7). Comets C/2003 K10, C/2003 L5, C/2003 M1, C/2003 M2, and C/2003 M5 were also visible in the C3 coronagraph. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 K10 May 26.571  4^{h}14.^{m}8  +19^{o}00′  2004-G19
C/2003 K11  27.201  4 16.4  +19 39  2004-G19
C/2003 K12  29.118  4 20.5  +19 41  2004-G19
C/2003 L3  June 2.479  4 40.4  +20 29  2004-G19
C/2003 L4  5.814  4 53.3  +20 50  2004-G19
C/2003 L5  8.929  5 08.5  +18 33  2004-G20
C/2003 L6  9.814  5 08.9  +21 13  2004-G20
C/2003 L7  13.963  5 24.4  +21 25  2004-G20
C/2003 L8  15.621  5 31.3  +21 28  2004-H20
C/2003 M1  16.154  5 33.4  +21 24  2004-H20
C/2003 M2  17.721  5 39.9  +19 51  2004-H20
C/2003 M3  17.896  5 40.1  +21 33  2004-H20
C/2003 M4  18.479  5 40.4  +21 44  2004-H21
C/2003 M5  19.071  5 42.5  +20 01  2004-H21
C/2003 M6  19.746  5 45.6  +22 00  2004-H21
C/2003 M7  20.504  5 50.3  +21 42  2004-H21
2004 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green